Find the slope of the line through (-2,5) and (1,-7)

answer -4
why is the answer -4 need a full break down

Found 2 solutions by Mathtut, RAY100:
Answer by Mathtut(3670)

slope is rise over the run or the change in y over the change in x
:
(5-(-7))/(-2-1)=12/-3=-4 or (-7-5)/(1-(-2))=-12/3=-4
:
direction is important if you start with the first y coordinate for the change in y then you must also start with the first x coordinate for the change in x
and if you start with the 2nd y coorinate for the change in y then you MUST start with the 2nd x coordinate for the change in x.
:

Answer by RAY100(1637)

very good (-4) is slope
a simple sketch usually shows rise / run easily
just note end points, roughly position points and line, and calc x & y distances
to calculate, use m=(y2-y1) / (x2-x1)= ( -7 -5) / 1-(-2)) = -12/3=-4